OpenStudy (anonymous):

Derivative of Sqrt(x) - 3 / Sqrt (x) + 3 Quotient rule, or another way? If you don't mind explaining how you got your answer, that would be awesome! Derivative of Sqrt(x) - 3 / Sqrt (x) + 3 Quotient rule, or another way? If you don't mind explaining how you got your answer, that would be awesome! @Mathematics

OpenStudy (agreene):

\[\frac{d}{dx} \frac{\sqrt{x}-3}{\sqrt{x}+3}\] is that what you are asking? If so, Quotient rule would probably be your best bet.

OpenStudy (anonymous):

When i do it, the top gets wiped out ...

OpenStudy (agreene):

your final answer is \[\frac{3}{(\sqrt{x}+3)^2\sqrt{x}}\]

OpenStudy (anonymous):

I will write it out , so If I made an error it can be pointed out. \[((x^\left\{ .5 \right\}-3)(.5x^\left\{ -3/2 \right\})-(x^\left\{.5 \right\}+3)(.5x^\left\{ -3/2 \right\}))/(\sqrt{x}+3)^2 \]

OpenStudy (anonymous):

\[(0.5x^{-1}-1.5x^{-3/2}-0.5x^{-1}+1.5x^{-3/2})/(\sqrt{x}+3)\]

OpenStudy (anonymous):

^2*

OpenStudy (anonymous):

the answer you provided wasn't accepted by the web work.

OpenStudy (anonymous):

never mind it was the right answer, I forgot to plug in a value .... HUman error sorry. But can you show how you got it, if its not to much to ask ! :-)

OpenStudy (agreene):

yeah, I'm not sure what you are doing. Quotient rule: \[\frac{d}{dx}\frac{u}{v}=\frac{v\frac{du}{dx}-u\frac{du}{dx}}{v^2}\] where: u=x^(1/2)-3 and v=x^(1/2)+3 which leads us to: \[\frac{(\sqrt{x}-3)(\frac{d}{dx}(\sqrt{x}-3))-(\sqrt{x}+3)(\frac{d}{dx}(\sqrt{x}+3))}{(\sqrt{x}+3)^2}\] as the first step.

OpenStudy (agreene):

\[\frac{(\sqrt x-3)(\frac{d}{dx}(-3)+\frac{d}{dx}(\sqrt x))-(\sqrt x +3)(\frac{d}{dx}(3)+\frac{d}{dx}(\sqrt x)) }{(\sqrt{x}+3)^2}\] now that everything is separated we can start doing derivatives :D \[\frac{(\sqrt x-3)(\frac{1}{2\sqrt x})+(\sqrt x+3)(\frac{1}{2 \sqrt x})}{(\sqrt x +3)^2}\] simplifying a touch: \[\large \frac{\frac{\sqrt x +3}{2 \sqrt x}-\frac{\sqrt x-3}{2\sqrt x }}{(\sqrt x +3 )^2}\] from there you can evaluate and get to the answer I first said :D

OpenStudy (anonymous):

thanks man